A steel beam 12.0m sits next to a concrete wall when the temperature is 20.0°C.

A gap must be left between the beam and the concrete wall for expansion

purposes. If the temperature rises to 45.0°C, how large must the gap be if the

steel beam just touches the concrete wall?


1
Expert's answer
2021-09-24T09:26:13-0400

Given:

"l_0=12.0\: \rm m"

"\alpha=12.0*10^{-6}\: \rm K^{-1}"

"\Delta t=25\:\rm K"


The size of a gap

"\Delta l=l_0*\alpha\Delta t\\=12.0\:\rm m*12.0*10^{-6}\: K^{-1}*25\: K=0.0036\: m=3.6\: mm"